When ordering new desks for classrooms, the business manager for a college uses the fact that there are approx. 9 times as many right-handed students as left handed. how many new desks of ecah type must be ordered if 600 new desks are needed?
How do I set this up?

Answer by checkley77(12844)

9n+1n=600
10n=600
n=600/10
n=60 left handed desks.
9*n=540 right handed desks.